OpenStudy (anonymous):

The graph of the equation given below is a circle. What is the length of the radius of the circle? (x - 6)2 + (y + 5)2 = 212

OpenStudy (anonymous):

rewrite this to (x-6)+(y+5)=106

OpenStudy (mertsj):

\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (mertsj):

That is the equation of a circle. Center is (h,k) and radius = r

OpenStudy (anonymous):

yes Im sorry forgot the power of twos

OpenStudy (mertsj):

Got the center right, but NOT the radius

OpenStudy (anonymous):

im sorry i meant to write ** =21^2